OpenStudy (unklerhaukus):

Disjunctions (a)\[(\pi>3)\vee(\pi>10)\] (b)\[(x<0)\vee(x>0)\] (c)\[(x=0)\vee(x>0)\] (d)\[(x>0)\vee(x\geq0)\] (e)\[(x>3)\vee(x^2>9)\]

OpenStudy (unklerhaukus):

(a) \[(\pi>3)\vee(\pi>10)\quad\longrightarrow\quad \pi>10\] (b) \[(x<0)\vee(x>0)\quad\longrightarrow\quad x\neq0\] (c) \[(x=0)\vee(x>0)\quad\longrightarrow\quad x\geq 0\] (d) \[(x>0)\vee(x\geq0)\quad\longrightarrow\quad x\geq0\] (e) \[(x>3)\vee(x^2>9)\quad\longrightarrow\quad (x>3)\vee[x>3 \vee x<-3]\]\[\qquad\qquad\qquad\qquad\qquad\longrightarrow\quad (x>3)\vee (x<-3)\]\[\qquad\qquad\qquad\qquad\qquad\longrightarrow\quad (x^2>9)\]

OpenStudy (unklerhaukus):

please check my work

hartnn (hartnn):

how (a) ?? pi>3 ?

OpenStudy (unklerhaukus):

hmm yeah

hartnn (hartnn):

rest are correct.

OpenStudy (unklerhaukus):

thank you !